EXAMINATION OF H_{2}CO–X^{+} AND NH_{2}CH_{2}COOH–X^{+} COMPLEXES [X^{+} = Li^{+}, Na^{+}, K^{+}] USING ELEC-TRONIC STRUCTURE THEORY

\underline{ANNIE L. LESIAK}, SAMANTHA HORVATH, and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Infrared spectroscopy is a powerful tool for studying molecular structure as changes in the frequency and intensity of infrared transitions provide a way to probe environment effects on molecular systems. In this study we investigate how the frequency and intensity of the C=O stretch vibration of formaldehyde and glycine change upon the introduction of alkali metal cations. Specifically we focus on the complexes of Li^{+}, Na^{+}, and K^{+} with H_{2}CO and NH_{2}CH_{2}COOH. There is evidence of small changes in the harmonic C=O stretch frequency; however, the changes in intensity are much larger. For example the intensity of the C=O stretch in H_{2}CO–Li^{+} is twice that in bare H_{2}CO. The ultimate goal of this work is to determine how the sensitivity of the C=O stretch vibration is affected by the particular alkali metal cation as well as the geometry of the complex. All calculations in this study were performed using Gaussian03 at the MP2/6-311G^{∗} level of theory/basis set.